i am trying but i am getting frustrated!!! what is the greatest common factor of 10280 and 6988? are we able to find all the common factors. if not is there an easier way for me to find this? Please help!!!
Found 2 solutions by Alan3354, scott8148:
Answer by Alan3354(30993)

10280 = 4*10*257
---------
6988 = 4*1747
The GCF is 4

Answer by scott8148(6628)

factor using PRIME numbers, starting with 2 and working up

10280/2 = 5140
/2 = 2570
/2 = 1285 ___ not divisible by 3
/5 = 257 ___ this is prime
10280 = 2^3 * 5 * 257

6988/2 = 3494
/2 = 1747 ___ this is prime
6988 = 2^2 * 1747

so the GCF is 2^2